# Computational Complexity of Physical Information Sufficiency

### v3

@@ A.1 Lean Handle ID Map
CC41 DecisionQuotient.ClaimClosure.no_uncertified_exact_claim_core
CC42 DecisionQuotient.ClaimClosure.one_step_bridge
CC43 DecisionQuotient.ClaimClosure.oracle_lattice_transfer_as_regime_simulation
CC44 DecisionQuotient.ClaimClosure.physical_crossover_above_cap_core
CC45 DecisionQuotient.ClaimClosure.physical_crossover_core
CC46 DecisionQuotient.ClaimClosure.physical_crossover_hardness_core
CC47 DecisionQuotient.ClaimClosure.physical_crossover_policy_core
CC48 DecisionQuotient.ClaimClosure.process_bridge_failure_witness
CC49 DecisionQuotient.ClaimClosure.query_obstruction_boolean_corollary
CC50 DecisionQuotient.ClaimClosure.query_obstruction_finite_state_core
CC51 DecisionQuotient.ClaimClosure.regime_core_claim_proved
CC52 DecisionQuotient.ClaimClosure.regime_simulation_transfers_hardness
CC53 DecisionQuotient.ClaimClosure.reusable_heuristic_of_detectable
CC54 DecisionQuotient.ClaimClosure.selectorSufficient_not_implies_setSufficient
CC55 DecisionQuotient.ClaimClosure.separable_detectable
CC56 DecisionQuotient.ClaimClosure.snapshot_vs_process_typed_boundary
CC57 DecisionQuotient.ClaimClosure.standard_assumption_ledger_unpack
CC58 DecisionQuotient.ClaimClosure.stochastic_objective_bridge_can_fail_on_sufficiency
CC59 DecisionQuotient.ClaimClosure.subproblem_hardness_lifts_to_full
CC60 DecisionQuotient.ClaimClosure.subproblem_transfer_as_regime_simulation
CC61 DecisionQuotient.ClaimClosure.sufficiency_conp_complete_conditional
CC62 DecisionQuotient.ClaimClosure.sufficiency_conp_reduction_core
CC63 DecisionQuotient.ClaimClosure.sufficiency_iff_dq_ratio
CC64 DecisionQuotient.ClaimClosure.sufficiency_iff_projectedOptCover_eq_opt
CC65 DecisionQuotient.ClaimClosure.thermo_conservation_additive_core
CC66 DecisionQuotient.ClaimClosure.thermo_energy_carbon_lift_core
CC67 DecisionQuotient.ClaimClosure.thermo_eventual_lift_core
CC68 DecisionQuotient.ClaimClosure.thermo_hardness_bundle_core
CC69 DecisionQuotient.ClaimClosure.thermo_mandatory_cost_core
CC70 DecisionQuotient.ClaimClosure.tractable_bounded_core
CC71 DecisionQuotient.ClaimClosure.tractable_separable_core
CC72 DecisionQuotient.ClaimClosure.tractable_subcases_conditional
CC73 DecisionQuotient.ClaimClosure.tractable_tree_core
CC74 DecisionQuotient.ClaimClosure.transition_coupled_bridge_can_fail_on_sufficiency
CC75 DecisionQuotient.ClaimClosure.tree_structure_detectable
CC76 DecisionQuotient.ClaimClosure.typed_claim_admissibility_core
CC77 DecisionQuotient.ClaimClosure.typed_static_class_completeness
@@
CT1 CT.law_gap_physical_claim_holds
CT2 CT.no_law_gap_counterexample
CT3 CT.no_physical_counterexample_of_core_theorem
CT4 CT.physical_bridge_bundle
@@
CT5 CT.physical_claim_lifts_from_core
@@
CT6 CT.physical_claim_lifts_from_core_conditional
@@
CT7 CT.physical_counterexample_invalidates_core_rule
CT8 CT.physical_counterexample_yields_core_counterexample
@@
DP6 DecisionProblem.anchorSufficient
DP7 DecisionProblem.classMonotoneOn
DP8 DecisionProblem.constant_opt_all_sufficient
DP21 DecisionProblem.edgeOnComplement_iff_not_sufficient
DQ3 DecisionQuotient.card_anchoredSlice_eq_pow_sub
DQ5 DecisionQuotient.opt_eq_feasible_of_gap
DQ6 DecisionQuotient.opt_eq_of_allowed_iff
DQ7 DecisionQuotient.opt_singleton_of_logicallyDeterministic
DQ8 DecisionQuotient.universe_sets_objective_schema
DQ154 DecisionQuotient.Physics.Instantiation.DecisionCircuit
DQ155 DecisionQuotient.Physics.Instantiation.DecisionInterpretation
DQ159 DecisionQuotient.Physics.Instantiation.MoleculeAsCircuit
DQ160 DecisionQuotient.Physics.Instantiation.MoleculeAsDecisionCircuit
DQ517 DecisionQuotient.Physics.Instantiation.Circuit
DQ518 DecisionQuotient.Physics.Instantiation.Dynamics
DQ519 DecisionQuotient.Physics.Instantiation.Geometry
DQ533 DecisionQuotient.Physics.DecisionTime.decisionTrace_length_eq_ticks
DQ534 DecisionQuotient.Physics.DecisionTime.decision_count_equals_elapsed_time
DQ535 DecisionQuotient.Physics.DecisionTime.decision_event_iff_eq_tick
@@
DQ537 DecisionQuotient.Physics.DecisionTime.decision_taking_place_is_unit_of_time
DQ539 DecisionQuotient.Physics.DecisionTime.run_elapsed_time_eq_ticks
DQ540 DecisionQuotient.Physics.DecisionTime.run_time_exact
@@
DQ541 DecisionQuotient.Physics.DecisionTime.substrate_step_is_time_unit
DQ542 DecisionQuotient.Physics.DecisionTime.substrate_step_realizes_decision_event
DQ545 DecisionQuotient.Physics.DecisionTime.tick_increments_time
DQ546 DecisionQuotient.Physics.DecisionTime.tick_is_decision_event
DQ547 DecisionQuotient.Physics.DecisionTime.time_coordinate_falsifiable
@@
DQ548 DecisionQuotient.Physics.DecisionTime.time_is_discrete
DQ549 DecisionQuotient.Physics.DecisionTime.time_unit_law_substrate_invariant
DQ560 DecisionQuotient.HardnessDistribution.hardnessLowerBound
DQ561 DecisionQuotient.HardnessDistribution.hardness_is_irreducible_required_work
DQ562 DecisionQuotient.HardnessDistribution.requiredWork
DQ563 DecisionQuotient.HardnessDistribution.requiredWork_eq_affine_in_sites
DQ564 DecisionQuotient.HardnessDistribution.workGrowthDegree
DQ565 DecisionQuotient.HardnessDistribution.workGrowthDegree_zero_iff_eventually_constant
DQ566 DecisionQuotient.Physics.PhysicalCore.admissible_report_requires_strict_certified_work
DQ567 DecisionQuotient.Physics.PhysicalCore.certified_work_lower_bound_raw
DQ568 DecisionQuotient.Physics.PhysicalCore.finite_budget_blocks_canonical_exponential_certification
DQ569 DecisionQuotient.Physics.PhysicalCore.finite_budget_blocks_exact_certification_beyond_threshold
DQ570 DecisionQuotient.Physics.PhysicalCore.rlff_abstain_strict_global_maximizer_of_no_certificates
DQ571 DecisionQuotient.Physics.PhysicalCore.substrate_transport_blocks_encoded_counterexamples
DQ572 DecisionQuotient.Physics.PhysicalCore.substrate_transport_sound_conditional
DQ573 DecisionQuotient.Physics.PhysicalCore.eventual_no_exact_admissibility_under_budget
DQ574 DecisionQuotient.Physics.PhysicalCore.eventual_no_exact_admissibility_under_canonical_requirement
DQ575 DecisionQuotient.PhysicalBudgetCrossover.crossover_eventually_of_eventual_split
DQ576 DecisionQuotient.PhysicalBudgetCrossover.exists_least_crossover_point
DQ577 DecisionQuotient.PhysicalBudgetCrossover.explicit_eventual_infeasibility_of_monotone_and_witness
DQ578 DecisionQuotient.PhysicalBudgetCrossover.no_universal_survivor_without_succinct_bound
DQ579 DecisionQuotient.PhysicalBudgetCrossover.payoff_threshold_explicit_vs_succinct
DQ580 DecisionQuotient.PhysicalBudgetCrossover.policy_closure_at_divergence
DQ581 DecisionQuotient.PhysicalBudgetCrossover.policy_closure_beyond_divergence
DQ582 DecisionQuotient.Physics.PhysicalIncompleteness.under_resolution_implies_collision
DQ583 DecisionQuotient.Physics.PhysicalIncompleteness.under_resolution_implies_decision_collision
DQ584 DecisionQuotient.Physics.HeisenbergStrong.HeisenbergStrongBinding
DQ585 DecisionQuotient.Physics.HeisenbergStrong.strong_binding_implies_core_nontrivial
DQ586 DecisionQuotient.Physics.HeisenbergStrong.strong_binding_implies_nontrivial_opt_via_uncertainty
DQ587 DecisionQuotient.checking_witnessing_duality
DQ588 DecisionQuotient.checking_witnessing_duality_budget
@@
HD17 DecisionQuotient.HardnessDistribution.right_dominates_wrong
HD18 DecisionQuotient.HardnessDistribution.saturatingSiteCost_eventually_constant
HD19 DecisionQuotient.HardnessDistribution.simplicityTax_grows
HD20 DecisionQuotient.HardnessDistribution.totalDOF_eventually_constant_iff_zero_distributed
HD21 DecisionQuotient.HardnessDistribution.totalDOF_ge_intrinsic

### v4

@@ A.1 Lean Handle ID Map
CC54DecisionQuotient.ClaimClosure.posed_anchor_exact_claim_admissible_iff_competent
CC55DecisionQuotient.ClaimClosure.posed_anchor_exact_claim_requires_evidence
CC56DecisionQuotient.ClaimClosure.posed_anchor_no_competence_no_exact_claim
CC57DecisionQuotient.ClaimClosure.posed_anchor_query_truth_iff_exists_anchor
CC58DecisionQuotient.ClaimClosure.posed_anchor_query_truth_iff_exists_forall
CC59DecisionQuotient.ClaimClosure.posed_anchor_signal_positive_certified_implies_admissible
CC60DecisionQuotient.ClaimClosure.query_obstruction_boolean_corollary
CC61DecisionQuotient.ClaimClosure.query_obstruction_finite_state_core
CC62DecisionQuotient.ClaimClosure.regime_core_claim_proved
CC63DecisionQuotient.ClaimClosure.regime_simulation_transfers_hardness
CC64DecisionQuotient.ClaimClosure.reusable_heuristic_of_detectable
CC65DecisionQuotient.ClaimClosure.selectorSufficient_not_implies_setSufficient
CC66DecisionQuotient.ClaimClosure.separable_detectable
CC67DecisionQuotient.ClaimClosure.snapshot_vs_process_typed_boundary
CC68DecisionQuotient.ClaimClosure.standard_assumption_ledger_unpack
CC69DecisionQuotient.ClaimClosure.stochastic_objective_bridge_can_fail_on_sufficiency
CC70DecisionQuotient.ClaimClosure.subproblem_hardness_lifts_to_full
CC71DecisionQuotient.ClaimClosure.subproblem_transfer_as_regime_simulation
CC72DecisionQuotient.ClaimClosure.sufficiency_conp_complete_conditional
CC73DecisionQuotient.ClaimClosure.sufficiency_conp_reduction_core
CC74DecisionQuotient.ClaimClosure.sufficiency_iff_dq_ratio
CC75DecisionQuotient.ClaimClosure.sufficiency_iff_projectedOptCover_eq_opt
CC76DecisionQuotient.ClaimClosure.thermo_conservation_additive_core
CC77DecisionQuotient.ClaimClosure.thermo_energy_carbon_lift_core
CC78DecisionQuotient.ClaimClosure.thermo_eventual_lift_core
CC79DecisionQuotient.ClaimClosure.thermo_hardness_bundle_core
CC80DecisionQuotient.ClaimClosure.thermo_mandatory_cost_core
CC81DecisionQuotient.ClaimClosure.tractable_bounded_core
CC82DecisionQuotient.ClaimClosure.tractable_separable_core
CC83DecisionQuotient.ClaimClosure.tractable_subcases_conditional
CC84DecisionQuotient.ClaimClosure.tractable_tree_core
CC85DecisionQuotient.ClaimClosure.transition_coupled_bridge_can_fail_on_sufficiency
CC86DecisionQuotient.ClaimClosure.tree_structure_detectable
CC87DecisionQuotient.ClaimClosure.typed_claim_admissibility_core
CC88DecisionQuotient.ClaimClosure.typed_static_class_completeness
CC89DecisionQuotient.ClaimClosure.universal_solver_framing_core
CCC1DecisionQuotient.CC.anchor_sigma2p_complete_conditional
CCC2DecisionQuotient.CC.cost_asymmetry_eth_conditional
CCC3DecisionQuotient.CC.dichotomy_conditional
CCC4DecisionQuotient.CC.minsuff_collapse_to_conp_conditional
CCC5DecisionQuotient.CC.minsuff_conp_complete_conditional
CCC6DecisionQuotient.CC.sufficiency_conp_complete_conditional
CCC7DecisionQuotient.CC.tractable_subcases_conditional
CF1DecisionQuotient.Physics.ConstraintForcing.laws_not_determined_of_parameter_separation
CF2DecisionQuotient.Physics.ConstraintForcing.logic_time_not_sufficient_for_unique_law
CF3DecisionQuotient.Physics.ConstraintForcing.laws_determined_implies_objective_determined
CF4DecisionQuotient.Physics.ConstraintForcing.objective_not_determined_of_parameter_separation
CF5DecisionQuotient.Physics.ConstraintForcing.forcedDecisionBits_pos_of_deadline
CF6DecisionQuotient.Physics.ConstraintForcing.actionForced_of_deadline
CF7DecisionQuotient.Physics.ConstraintForcing.nondegenerateBelief_of_deadline_and_uncertainty
CF8DecisionQuotient.Physics.ConstraintForcing.forced_decision_implies_positive_landauer_cost
CF9DecisionQuotient.Physics.ConstraintForcing.forced_decision_implies_positive_nv_work
CH1DecisionQuotient.ClaimClosure.CH1
CH2DecisionQuotient.ClaimClosure.CH2
CH3DecisionQuotient.ClaimClosure.CH3
CH5DecisionQuotient.ClaimClosure.CH5
CH6DecisionQuotient.ClaimClosure.CH6
@@
CT1DecisionQuotient.Physics.ClaimTransport.PhysicalEncoding
@@
CT2DecisionQuotient.Physics.ClaimTransport.physical_claim_lifts_from_core
@@
CT3DecisionQuotient.Physics.ClaimTransport.physical_claim_lifts_from_core_conditional
CT4DecisionQuotient.Physics.ClaimTransport.physical_counterexample_yields_core_counterexample
@@
CT5DecisionQuotient.Physics.ClaimTransport.physical_counterexample_invalidates_core_rule
CT6DecisionQuotient.Physics.ClaimTransport.no_physical_counterexample_of_core_theorem
CT7DecisionQuotient.Physics.ClaimTransport.LawGapInstance
CT8DecisionQuotient.Physics.ClaimTransport.lawGapEncoding
CT9DecisionQuotient.Physics.ClaimTransport.lawGapPhysicalClaim
CT10DecisionQuotient.Physics.ClaimTransport.law_gap_physical_claim_holds
CT11DecisionQuotient.Physics.ClaimTransport.no_law_gap_counterexample
CT12DecisionQuotient.Physics.ClaimTransport.physical_bridge_bundle
CV1DecisionQuotient.Physics.Conversation.RecurrentCircuit
CV2DecisionQuotient.Physics.Conversation.CoupledConversation
CV3DecisionQuotient.Physics.Conversation.JointState
CV4DecisionQuotient.Physics.Conversation.tick_uses_shared_node
CV5DecisionQuotient.Physics.Conversation.tick_shared_is_merged_emissions
CV6DecisionQuotient.Physics.Conversation.channel_projection_eq_iff_quantized_eq
CV7DecisionQuotient.Physics.Conversation.clamp_projection_eq_iff_same_clamped_bit
CV10DecisionQuotient.Physics.Conversation.explanationGap_add_explanationBits
CV11DecisionQuotient.Physics.Conversation.toClaimReport
CV12DecisionQuotient.Physics.Conversation.abstain_iff_no_answer
CV13DecisionQuotient.Physics.Conversation.yes_no_iff_exact_claim
CV14DecisionQuotient.Physics.Conversation.toReportSignal_completion_defined_of_budget
CV15DecisionQuotient.Physics.Conversation.toReportSignal_signal_consistent_zero_certified
CV16DecisionQuotient.Physics.Conversation.abstain_report_can_carry_explanation
CV17DecisionQuotient.Physics.Conversation.clampDecisionEvent_iff_bitOps_pos
CV18DecisionQuotient.Physics.Conversation.clamp_event_implies_positive_energy
CV19DecisionQuotient.Physics.Conversation.BinaryAnswer
CV20DecisionQuotient.Physics.Conversation.ConversationReport
CV21DecisionQuotient.Physics.Conversation.explanationGap
CV26DecisionQuotient.Physics.Conversation.toClaimReport_ne_epsilon
CV27DecisionQuotient.Physics.Conversation.toReportSignal
CV28DecisionQuotient.Physics.Conversation.toReportSignal_declares_bound
DC1DecisionQuotient.StochasticSequential.static_stochastic_strict_separation
DC2DecisionQuotient.StochasticSequential.stochastic_sequential_strict_separation
DC3DecisionQuotient.StochasticSequential.complexity_dichotomy_hierarchy
DC4DecisionQuotient.StochasticSequential.regime_hierarchy
DC5DecisionQuotient.StochasticSequential.coNP_subset_PP
DC6DecisionQuotient.StochasticSequential.PP_subset_PSPACE
DC7DecisionQuotient.StochasticSequential.coNP_subset_PSPACE
DC8DecisionQuotient.StochasticSequential.static_to_coNP
DC9DecisionQuotient.StochasticSequential.stochastic_to_PP
DC10DecisionQuotient.StochasticSequential.sequential_to_PSPACE
DC11DecisionQuotient.StochasticSequential.ClaimClosure.claim_six_subcases
DC12DecisionQuotient.StochasticSequential.ClaimClosure.claim_hierarchy
DC13DecisionQuotient.StochasticSequential.ClaimClosure.claim_tractable_subcases_to_P
DC14DecisionQuotient.StochasticSequential.stochastic_dichotomy
DC15DecisionQuotient.StochasticSequential.above_threshold_hard
DC16DecisionQuotient.StochasticSequential.StochasticAnchorSufficient
DC17DecisionQuotient.StochasticSequential.StochasticAnchorSufficiencyCheck
DC18DecisionQuotient.StochasticSequential.stochastic_anchor_check_iff
DC19DecisionQuotient.StochasticSequential.stochastic_anchor_sufficient_of_stochastic_sufficient
DC20DecisionQuotient.StochasticSequential.SequentialAnchorSufficient
DC21DecisionQuotient.StochasticSequential.SequentialAnchorSufficiencyCheck
DC22DecisionQuotient.StochasticSequential.sequential_anchor_check_iff
DC23DecisionQuotient.StochasticSequential.sequential_anchor_sufficient_of_sequential_sufficient
DC24DecisionQuotient.StochasticSequential.StochasticAnchorCheckInstance
DC25DecisionQuotient.StochasticSequential.reduceMAJSAT_correct_anchor_strict
DC26DecisionQuotient.StochasticSequential.reduceMAJSAT_to_stochastic_anchor_reduction
DC27DecisionQuotient.StochasticSequential.SequentialAnchorCheckInstance
DC28DecisionQuotient.StochasticSequential.reduceTQBF_correct_anchor
DC29DecisionQuotient.StochasticSequential.reduceTQBF_to_sequential_anchor_reduction
DC30DecisionQuotient.StochasticSequential.StatePotential
DC31DecisionQuotient.StochasticSequential.utilityFromPotentialDrop_le_iff_nextPotential_ge
DC32DecisionQuotient.StochasticSequential.utility_from_action_state_potential
DC33DecisionQuotient.StochasticSequential.stochasticExpectedUtility_eq_neg_expectedActionPotential
DC34DecisionQuotient.StochasticSequential.stochasticExpectedUtility_le_iff_expectedActionPotential_ge
DC35DecisionQuotient.StochasticSequential.landauerEnergyFloor_nonneg
DC36DecisionQuotient.StochasticSequential.landauerEnergyFloor_mono_bits
DC37DecisionQuotient.StochasticSequential.thermodynamicCost_eq_landauerEnergyFloorRoom_states
DE1DecisionQuotient.ClaimClosure.DE1
DE2DecisionQuotient.ClaimClosure.DE2
DE3DecisionQuotient.ClaimClosure.DE3
DE4DecisionQuotient.ClaimClosure.DE4
DG1DecisionQuotient.Outside
DG2DecisionQuotient.anchoredSlice
DG3DecisionQuotient.anchoredSliceEquivOutside
DG4DecisionQuotient.card_outside_eq_sub
DG5DecisionQuotient.card_anchoredSlice
DG6DecisionQuotient.card_anchoredSlice_eq_pow_sub
DG7DecisionQuotient.card_anchoredSlice_eq_uniform
DG8DecisionQuotient.anchoredSlice_mul_fixed_eq_full
DG9DecisionQuotient.constantBoolDP
DG10DecisionQuotient.firstCoordDP
DG11DecisionQuotient.constantBoolDP_opt
DG12DecisionQuotient.firstCoordDP_opt
DG13DecisionQuotient.constantBoolDP_empty_sufficient
DG14DecisionQuotient.firstCoordDP_empty_not_sufficient
DG15DecisionQuotient.boolHypercube_node_count
DG16DecisionQuotient.node_count_does_not_determine_edge_geometry
DG17DecisionQuotient.DecisionProblem.edgeOnComplement
DG18DecisionQuotient.DecisionProblem.edgeOnComplement_iff_not_sufficient
@@
DP6DecisionQuotient.ClaimClosure.DP6
DP7DecisionQuotient.ClaimClosure.DP7
DP8DecisionQuotient.ClaimClosure.DP8
DQ1DecisionQuotient.ClaimClosure.DQ1
DQ2DecisionQuotient.ClaimClosure.DQ2
DQ3DecisionQuotient.ClaimClosure.DQ3
DQ4DecisionQuotient.ClaimClosure.DQ4
DQ5DecisionQuotient.ClaimClosure.DQ5
DQ6DecisionQuotient.ClaimClosure.DQ6
DQ7DecisionQuotient.ClaimClosure.DQ7
DQ8DecisionQuotient.ClaimClosure.DQ8
DS1DecisionQuotient.ClaimClosure.DS1
DS2DecisionQuotient.ClaimClosure.DS2
DS3DecisionQuotient.ClaimClosure.DS3
DS4DecisionQuotient.ClaimClosure.DS4
DS5DecisionQuotient.ClaimClosure.DS5
DS6DecisionQuotient.ClaimClosure.DS6
DT1DecisionQuotient.Physics.DecisionTime.TimedState
DT2DecisionQuotient.Physics.DecisionTime.DecisionProcess
DT3DecisionQuotient.Physics.DecisionTime.tick
DT4DecisionQuotient.Physics.DecisionTime.DecisionEvent
DT5DecisionQuotient.Physics.DecisionTime.TimeUnitStep
DT6DecisionQuotient.Physics.DecisionTime.time_is_discrete
DT7DecisionQuotient.Physics.DecisionTime.time_coordinate_falsifiable
DT8DecisionQuotient.Physics.DecisionTime.tick_increments_time
DT9DecisionQuotient.Physics.DecisionTime.tick_decision_witness
DT10DecisionQuotient.Physics.DecisionTime.tick_is_decision_event
@@
DT11DecisionQuotient.Physics.DecisionTime.decision_event_implies_time_unit
@@
DT12DecisionQuotient.Physics.DecisionTime.decision_taking_place_is_unit_of_time
DT13DecisionQuotient.Physics.DecisionTime.decision_event_iff_eq_tick
DT14DecisionQuotient.Physics.DecisionTime.run
DT15DecisionQuotient.Physics.DecisionTime.run_time_exact
DT16DecisionQuotient.Physics.DecisionTime.run_elapsed_time_eq_ticks
DT17DecisionQuotient.Physics.DecisionTime.decisionTrace
DT18DecisionQuotient.Physics.DecisionTime.decisionTrace_length_eq_ticks
@@
DT19DecisionQuotient.Physics.DecisionTime.decision_count_equals_elapsed_time
DT20DecisionQuotient.Physics.DecisionTime.SubstrateKind
DT21DecisionQuotient.Physics.DecisionTime.SubstrateModel
DT22DecisionQuotient.Physics.DecisionTime.substrate_step_realizes_decision_event
DT23DecisionQuotient.Physics.DecisionTime.substrate_step_is_time_unit
DT24DecisionQuotient.Physics.DecisionTime.time_unit_law_substrate_invariant
EI1DecisionQuotient.ThermodynamicLift.energy_ge_kbt_nat_entropy
EI2DecisionQuotient.ThermodynamicLift.energy_ground_state_tracks_entropy
EI4DecisionQuotient.ThermodynamicLift.landauerJoulesPerBit_pos
EI5DecisionQuotient.ThermodynamicLift.neukart_vinokur_duality
EP1DecisionQuotient.Physics.LocalityPhysics.landauer_principle
EP2DecisionQuotient.Physics.LocalityPhysics.finite_regional_energy
EP3DecisionQuotient.Physics.LocalityPhysics.finite_signal_speed
EP4DecisionQuotient.Physics.LocalityPhysics.nontrivial_physics
FI3DecisionQuotient.FunctionalInformation.functional_information_from_counting
FI6DecisionQuotient.FunctionalInformation.functional_information_from_thermodynamics
FI7DecisionQuotient.FunctionalInformation.first_principles_thermo_coincide
FN7DecisionQuotient.BayesOptimalityProof.KL_nonneg
FN8DecisionQuotient.BayesOptimalityProof.entropy_le_crossEntropy
FN12DecisionQuotient.BayesOptimalityProof.crossEntropy_eq_entropy_add_KL
FN14DecisionQuotient.BayesOptimalityProof.bayes_is_optimal
FN15DecisionQuotient.BayesOptimalityProof.KL_eq_sum_klFun
FN16DecisionQuotient.BayesOptimalityProof.KL_eq_zero_iff_eq
FP1DecisionQuotient.Physics.LocalityPhysics.trivial_states_all_equal
FP2DecisionQuotient.Physics.LocalityPhysics.equal_states_constant_function
FP3DecisionQuotient.Physics.LocalityPhysics.constant_function_singleton_image
FP4DecisionQuotient.Physics.LocalityPhysics.singleton_image_zero_entropy
FP5DecisionQuotient.Physics.LocalityPhysics.zero_entropy_no_information
FP6DecisionQuotient.Physics.LocalityPhysics.triviality_implies_no_information
FP7DecisionQuotient.Physics.LocalityPhysics.information_requires_nontriviality
FP8DecisionQuotient.Physics.LocalityPhysics.atypical_states_rare
FP9DecisionQuotient.Physics.LocalityPhysics.random_misses_target
FP10DecisionQuotient.Physics.LocalityPhysics.errors_accumulate
FP11DecisionQuotient.Physics.LocalityPhysics.wrong_paths_dominate
FP12DecisionQuotient.Physics.LocalityPhysics.second_law_from_counting
FP13DecisionQuotient.Physics.LocalityPhysics.verification_is_information
FP14DecisionQuotient.Physics.LocalityPhysics.entropy_is_information
FP15DecisionQuotient.Physics.LocalityPhysics.landauer_structure
FPT1DecisionQuotient.Physics.LocalityPhysics.Timeline
FPT2DecisionQuotient.Physics.LocalityPhysics.FPT2_function_deterministic
FPT3DecisionQuotient.Physics.LocalityPhysics.FPT3_outputs_differ_inputs_differ
FPT4DecisionQuotient.Physics.LocalityPhysics.FPT4_step_requires_distinct_moments
FPT5DecisionQuotient.Physics.LocalityPhysics.FPT5_distinct_moments_positive_duration
FPT6DecisionQuotient.Physics.LocalityPhysics.FPT6_step_takes_positive_time
FPT7DecisionQuotient.Physics.LocalityPhysics.FPT7_no_instantaneous_steps
FPT8DecisionQuotient.Physics.LocalityPhysics.FPT8_propagation_takes_time
FPT9DecisionQuotient.Physics.LocalityPhysics.FPT9_speed_bounded_by_positive_time
FPT10DecisionQuotient.Physics.LocalityPhysics.FPT10_ec3_is_logical
FS1DecisionQuotient.Statistics.sum_fisherScore_eq_srank
FS2DecisionQuotient.Statistics.fisherMatrix_rank_eq_srank
FS3DecisionQuotient.Statistics.fisherMatrix_trace_eq_srank
FS4DecisionQuotient.Statistics.fisherScore_relevant
FS5DecisionQuotient.Statistics.fisherScore_irrelevant
GE1DecisionQuotient.ClaimClosure.GE1
GE2DecisionQuotient.ClaimClosure.GE2
GE3DecisionQuotient.ClaimClosure.GE3
GE4DecisionQuotient.ClaimClosure.GE4
GE5DecisionQuotient.ClaimClosure.GE5
GE6DecisionQuotient.ClaimClosure.GE6
GE7DecisionQuotient.ClaimClosure.GE7
GE8DecisionQuotient.ClaimClosure.GE8
GE9DecisionQuotient.ClaimClosure.GE9
GN1DecisionQuotient.LogicGraph.isCycle
GN2DecisionQuotient.LogicGraph.cycleWitnessBits_pos
GN3DecisionQuotient.LogicGraph.pathProb_nonneg
GN4DecisionQuotient.LogicGraph.pathSurprisal_nonneg_of_positive_mass
GN5DecisionQuotient.LogicGraph.nontrivialityScore_unknown
GN6DecisionQuotient.LogicGraph.observerEntropy_nonneg
GN7DecisionQuotient.LogicGraph.dqFromEntropy_in_unit_interval
GN8DecisionQuotient.LogicGraph.path_belief_forced_under_uncertainty
GN9DecisionQuotient.LogicGraph.bayes_update_exists_for_observer_paths
GN10DecisionQuotient.LogicGraph.cycle_witness_implies_positive_landauer
GN11DecisionQuotient.LogicGraph.cycle_witness_implies_positive_nv_work
GN12DecisionQuotient.LogicGraph.dna_erasure_implies_positive_landauer
GN13DecisionQuotient.LogicGraph.dna_room_temp_environmental_stability
@@
HD17DecisionQuotient.HardnessDistribution.requiredWork
HD18DecisionQuotient.HardnessDistribution.requiredWork_eq_affine_in_sites
HD19DecisionQuotient.HardnessDistribution.right_dominates_wrong
HD20DecisionQuotient.HardnessDistribution.saturatingSiteCost_eventually_constant
HD21DecisionQuotient.HardnessDistribution.simplicityTax_grows
HD22DecisionQuotient.HardnessDistribution.hardnessLowerBound
HD23DecisionQuotient.HardnessDistribution.hardness_is_irreducible_required_work
HD24DecisionQuotient.HardnessDistribution.totalDOF_eventually_constant_iff_zero_distributed
HD25DecisionQuotient.HardnessDistribution.totalDOF_ge_intrinsic
HD26DecisionQuotient.HardnessDistribution.totalExternalWork_eq_n_mul_gapCard
HD27DecisionQuotient.HardnessDistribution.workGrowthDegree
HD28DecisionQuotient.HardnessDistribution.workGrowthDegree_zero_iff_eventually_constant
HS1DecisionQuotient.Physics.HeisenbergStrong.NoisyPhysicalEncoding
HS2DecisionQuotient.Physics.HeisenbergStrong.HeisenbergStrongBinding
HS3DecisionQuotient.Physics.HeisenbergStrong.strong_binding_implies_core_nontrivial
HS4DecisionQuotient.Physics.HeisenbergStrong.strong_binding_yields_core_encoding_witness
HS5DecisionQuotient.Physics.HeisenbergStrong.strong_binding_implies_physical_nontrivial_opt_assumption
HS6DecisionQuotient.Physics.HeisenbergStrong.strong_binding_implies_nontrivial_opt_via_uncertainty
IA1DecisionQuotient.ClaimClosure.IA1
IA2DecisionQuotient.ClaimClosure.IA2
IA3DecisionQuotient.ClaimClosure.IA3
IA4DecisionQuotient.ClaimClosure.IA4
IA5DecisionQuotient.ClaimClosure.IA5
IA6DecisionQuotient.ClaimClosure.IA6
IA7DecisionQuotient.ClaimClosure.IA7
IA8DecisionQuotient.ClaimClosure.IA8
IA9DecisionQuotient.ClaimClosure.IA9
IA10DecisionQuotient.ClaimClosure.IA10
IA11DecisionQuotient.ClaimClosure.IA11
IA12DecisionQuotient.ClaimClosure.IA12
IA13DecisionQuotient.ClaimClosure.IA13